OpenStudy (anonymous):

If tan x = .67 and cos x = .83, what is sin x?

OpenStudy (anonymous):

\[\tan(x)=\frac{\sin(x)}{\cos(x)}\] \[\sin(x)=\cos(x)\times \tan(x)\]